### v1

@@ 2.2 Directed Acyclic Graph
To better extract information from quantum circuits, we can use Directed Acyclic Graphs (DAGs) to represent quantum gate circuits and apply them to tasks such as quantum architecture search [70, 72] and quantum compiling [107, 19]. By adding a source node and a sink node at the beginning and end of the quantum circuit, respectively, and treating each quantum gate as a node, we create a directed graph. Adjacent gates on the same qubit are connected by directed edges, and multi-qubit gates introduce additional edges. These directed edges represent the sequence in which the gates are applied. The advantages of using DAGs to represent quantum circuits are: (1) the structure of quantum circuits naturally transforms into DAGs; (2) DAGs can represent the application order of quantum gates and the control qubits of multi-qubit gates, maximizing the advantages of DAGs ; and (3) graph neural networks can effectively extract and compute features of nodes and edges, handling large-scale quantum circuits.
@@
Additionally, there is a special circuit representation in the form of directed acyclic graphs called Quantum Multiple-Valued Decision Diagram (QMDD) [143]. As illustrated in Figure 2, starting from the $q_{2}$ level, which is a $2\times 2$ matrix, we build the diagram layer by layer according to the transformation matrix. Each $2\times 2$ matrix consists of four matrix elements. There are four types of $q_{2}$ level matrices in the circuit, so this level has four nodes. The $q_{1}$ level above comprises a $4\times 4$ matrix. Each $q_{1}$ level matrix is composed of four $q_{2}$ level matrices, connecting each $q_{1}$ level node to four $q_{2}$ level nodes. The final $q_{0}$ level $8\times 8$ unitary matrix is the target unitary matrix we aim to build. The completed QMDD diagram is shown at the bottom of the QMDD module in Figure 2.
2.3 Circuit Polynomial
@@
Based on the representation of the quantum gate model, we can further characterize quantum circuits using algebraic methods, such as linear Boolean functions or phase polynomials [8, 136, 128, 6, 183]. These polynomial representation algorithms focus more on the quantum state rather than the specific quantum gates acting on the state. An arbitrary single-qubit quantum gate can be seen as inducing the following transformations on the quantum state:
|$\ket{x}\rightarrow e^{\text{i}p(x)}\ket{g(x)},$ (5)
where |$\ket{x}\rightarrow\ket{g(x)}$ is a reversible transformation, and $e^{\text{i}p(x)}$ represents the phase information imparted by the quantum gate to the current quantum state |$\ket{x}$. Since reversible transformations in quantum circuits are relatively straightforward to analyze, phase polynomial expressions focus more on the phase information contributed by quantum gates, particularly gates like Z, S, T, Rz$(\theta)$ and CZ. Given that the T gate is a crucial target in quantum circuit optimization, phase polynomial expressions play a significant role in optimizing circuits. Below, we present a specific example of a phase polynomial. For an $n$-qubit circuit, assuming that there are only two types of gates, CNOT and Rz$(\theta)$, the phase function $p(x)$ for the sample quantum circuit in Figure 2 can be expressed as:
$p(x_{1},x_{2},x_{3})=\theta_{1}(x_{1})+\theta_{2}(x_{1}\oplus x_{2})+\theta_{3}(x_{1}\oplus x_{2}\oplus x_{3})+\theta_{4}(x_{2}\oplus x_{3}).$ (6)
@@
Since the coefficients in the phase function are all rotation angles limited to a range of $2\pi$, we can take the modulo of the superimposed angle. This allows the phase function to serve as a criterion for determining whether two circuits are equivalent. Compared to directly comparing two unitary matrices, using phase polynomials enables the equivalence assessment of larger-scale circuits. For instance, Ruiz et al. [163] optimized a 72-qubit quantum circuit based on phase polynomial, a task that cannot be accomplished by merely comparing unitary matrices.
2.4 Tensor Networks
@@
As introduced by [151] for the mathematical concept of tensor networks, tensor can be regarded as a multidimensional vector, where scalars, vectors, and matrices can be considered as 0-dimensional, 1-dimensional, and 2-dimensional tensors, respectively [146]. Deutsch introduced the graphical representation of tensors into quantum computing, leading to the quantum gate circuits we discussed earlier. Therefore, quantum gate models are a special form of tensor networks. We will briefly introduce this more fundamental representation, with more detailed content about the tensor network available in [22]. Tensor networks can connect different tensors through indices, an operation called tensor network contraction. Some classic tensor network structures include Matrix Product State (MPS), Projected Entangled Pair States (PEPS), Tree Tensor Network (TTN), and Multi-Scale Entanglement Renormalization Ansatz (MERA), as shown in Figure 2.
@@
Since matrix multiplication can be viewed as the contraction of two second-order tensors based on the same index, quantum circuits can naturally be simulated and optimized using tensor networks [120]. The main optimization method is to reduce the cost of contractions, but finding an optimal tensor network contraction sequence is NP-hard [34]. This problem can be further modeled as finding an optimal tree decomposition of the tensor network graph structure [168]. The tree pruning method proposed in the paper can decompose the original graph into multiple parallel subgraphs with the minimum tree depth. Huang et al. [78] proposed index slicing, which divides the entire contraction into sub-tasks that can be executed in parallel. Specifically, a greedy algorithm selects indices that lead to the greatest reduction in complexity, accompanied by local reordering of the contraction tree. Additionally, pre-computing small tensors on the CPU and reordering matrix multiplication on the GPU can reduce the actual running cost. Contraction tasks can be divided into several tensor transpositions and one matrix multiplication [116]. Vincent et al. [185] constructed a task graph based on the dependencies of the aforementioned transpositions and multiplications, allowing different transpositions to be parallelized. Furthermore, by naming to ensure that shared operations between sliced contractions are only calculated once, costs are significantly reduced.
@@
Another research direction is to optimize performance by simulating various tensor network structures when designing quantum circuits. These circuits are often referred to as Tensor Network Quantum Circuits (TNQC). Generative and adversarial tasks are completed using quantum circuits based on MPS and TTN [84]. Since TNQC allows us to simulate certain quantum circuits efficiently, we can use this method to save physical qubits and reduce the impact of real quantum computing noise. Quantum circuits based on MPS and MERA are also analyzed [67], finding that TNQC can use fewer quantum gate parameters to achieve the same accuracy, further validating the practicality of TNQC.
2.5 ZX Diagrams
@@
ZX diagrams are a method for representing quantum circuits based on ZX calculus, proposed by Coecke and Duncan in 2008 [36]. ZX diagrams are also a special form of tensor networks that can represent any linear mapping between qubits [93]. They use spiders and the lines connecting spiders to represent a quantum circuit. Typically, ZX diagrams include two types of spiders: green Z spiders and red X spiders, as shown in Figure 2 [203]. When the angles are $\pi$, these two types of spiders correspond to the Pauli-Z and Pauli-X gates, respectively.
@@
Spiders define the basic linear operations in any dimension, where the lines on the left side of a spider represent the inputs of the current linear operation, and the lines on the right side represent the outputs. Each spider’s output can further connect to the inputs of other spiders, constructing a large network. ZX diagrams follow a set of fusion rules, which allow for obtaining many equivalent representations, simplifying the connections of complex spiders. Since 2019, much work has been done to optimize quantum circuits based on ZX diagrams. For instance, [49, 96, 51, 182] studied the optimization of the number of spiders in ZX diagrams; [44, 75, 134, 94] researched the optimization of T gate counts; and papers like [175, 59, 193, 92] introduced heuristic algorithms for the contraction and optimization of ZX diagrams. By optimizing the ZX diagram representation of quantum circuits, we can obtain more streamlined quantum circuits. When these ZX diagrams are mapped back to gate circuits, we can get circuits that meet the requirements. Compared to directly optimizing quantum gate circuits, the optimization methods based on ZX diagrams can introduce different characteristics, making them a relatively common method for representing quantum circuits.
3 Quantum Logic Circuit Synthesis
@@
In this section, we address the challenge of quantum logic circuit synthesis. Running a quantum algorithm on a quantum processor necessitates translating the algorithm’s unitary transformations into quantum circuits. Typically, these circuits are designed manually [87] or through decomposition algorithms [97, 184, 43]. This process demands substantial human labor and expertise, often resulting in circuits that are particularly deep. In the current Noisy Intermediate-Scale Quantum (NISQ) era, various constraints must be considered, including the accuracy of single and double qubit gates, qubit decoherence time, and the topological structure of physical qubits. Constructing a quantum circuit that achieves the highest possible accuracy while adhering to these constraints is a pressing challenge. Therefore, this section delves into methods for automatically generating quantum circuits that meet these requirements. This automated synthesis of quantum logic circuits is also referred to as the Quantum Architecture Search (QAS) problem in literature [209, 48, 115].
3.1 Problem Definition
@@
First, we provide the mathematical definition of the QAS problem. The QAS problem originates from the Neural Architecture Search (NAS) problem, which automates the synthesis and optimization of neural network structures and parameters to replace human expert design. NAS has been widely applied in the field of computer vision, such as image classification [111] and object detection [188], and has achieved success on multiple datasets. Similar to the NAS setting, when defining the QAS problem, we first determine the set of candidate gates and the search space. Assuming we need to search for a circuit with $n$ qubits and $m$ layers, define the set of candidate gates as $\mathcal{G}$, and the size of the candidate set as $k=|\mathcal{G}|$, then the search space of the QAS problem is of the size $k^{n\times m}$. The candidate gate set $\mathcal{G}$ can include single-qubit, two-qubit, and parameterized quantum gates. We denote the selected candidate gates as matrix $\mathbf{M}\in\mathcal{G}^{n\times m}$, define $\mathbf{U}{ij}=\sigma(\mathbf{M}{ij})$, where $\sigma$ maps a quantum gate on the $i$-th qubit and $j$-th layer to a $2^{n}\times 2^{n}$ unitary matrix with all the irrelevant qubits as identity. The unitary matrix corresponding to the quantum circuit obtained through the search can be expressed as:
@@
For different applications, the QAS problem can be further divided into two versions. The first one searches the circuit with a target unitary matrix $\mathbf{U}$ and the corresponding optimization objective can be expressed as:
$\min||\mathbf{U}-\hat{\mathbf{U}}||_{F},$ (9)
where $||\cdot||_{F}$ represents the F-norm of the matrix. The second one designs the circuit based on a series of quantum initial states |$\ket{\chi_{i}}$ and corresponding final states |$\ket{\psi_{i}}$, without an explicit unitary matrix. The corresponding optimization objective can be expressed as:
$\max\sum_{i}|\bra{\psi_{i}}\hat{\mathbf{U}}\ket{\chi_{i}}|^{2}.$ (10)
@@
3.2 Applications of Quantum Architecture Search
3.2.1 Quantum Architecture Search with Unitary Matrix
The most important application for QAS with unitary matrix is the circuit implementation of oracles. Here, an oracle refers to a black-box unitary matrix often seen in quantum algorithms. We may not know its specific implementation yet, but we understand the unitary operation corresponding to this oracle. Quantum oracle is of fundamental significance in quantum computing and quantum information, but implementing an arbitrary quantum oracle requires exponentially many gates [141]. Taking the oracle in Grover’s algorithm [64] as an example, we denote the oracle as $\mathcal{O}$, which can be represented as a unitary operation:
@@ 3.2.1 Quantum Architecture Search with Unitary Matrix
where |$\ket{x}$ is the register recording the index during the search, $\oplus$ is modulo $2$ addition, and |$\ket{q}$ is ancilla qubit. When $f(x)=1$, the ancilla qubit will flip; otherwise, no operation is performed. Therefore, in the searching process of Grover’s algorithm, we can prepare |$\ket{x}\ket{0}$, then apply the oracle $\mathcal{O}$ and determine whether $x$ is the solution to the search problem by checking if the ancilla qubit is |$\ket{1}$.
@@
Researchers typically manually design these oracles using basic quantum gates (i.e., quantum gates without parameters) [23, 158]. However, with the proposal of QAS methods, parameterized quantum circuits have also been used to search for implementations of these oracles more efficiently. Search algorithms are used to evolve quantum oracles [47] as well as design quantum adders [108, 45], which is further applied to quantum autoencoders[103]. Recent work [209] has also used QAS algorithms to design circuits for Quantum Fourier Transformations.
3.2.2 Quantum Architecture Search without Unitary Matrix
@@
In practical applications, situations without a target unitary matrix are more common, especially in quantum machine learning problems such as Quantum Neural Networks (QNN) or Variational Quantum Eigensolvers (VQE). Theoretically, if there are enough linearly independent input-output pairs (quantum initial states and corresponding final states), the unitary matrix corresponding to the transformation circuit is mathematically solvable, making this situation equivalent to the case with a target unitary matrix. However, the search process becomes more challenging when there is a lack of linearly independent input-output pairs.
@@
The VQE algorithm [152] is a notable application without a target unitary matrix, with extensive utility in quantum chemistry, combinatorial optimization, and quantum simulation. Its core functionality involves optimizing parameterized quantum circuits through classical optimizers to determine the minimum eigenvalue and corresponding eigenvector of a given Hamiltonian. VQE has garnered significant attention in the NISQ era due to its relatively shallow parameterized quantum circuits and adaptability to noisy environments on quantum processors. When utilizing VQE to solve quantum chemistry problems as the ground state energy estimation problem, the process typically involves discretizing electron distributions into orbitals through second quantization [21], mapping these orbitals to qubits using algorithms such as the Jordan-Wigner transformation [89], and employing unitary coupled cluster theory [179] to generate single and double excitation ansatzes for quantum state evolution. This procedure necessitates the derivation of distinct quantum circuits for each molecule or Hamiltonian, requiring both quantum chemistry expertise and specialized algorithms. Moreover, the resulting ansatzes are often deeper [17] due to the preservation of specific quantum chemical symmetries, rendering them challenging to implement on current NISQ hardware. Consequently, researchers have explored QAS methods to automatically design VQE ansatz, aiming to reduce design complexity and generate circuits that satisfy given constraints (e.g., circuit depth, two-qubit gate count, or physical qubit connectivity). The search process utilizes the same loss function as in the ground state energy estimation problem. Existing research has applied QAS methods to enhance unitary coupled cluster ansatzes [63, 165] or directly search for VQE circuits without prior ansatz knowledge [147, 187, 195]. Similarly, in combinatorial optimization problems, such as the maximum cut problem solved using the Quantum Approximate Optimization Algorithm (QAOA)[52], studies have proposed methods to optimize the original QAOA circuit[118, 119] or design related circuits that circumvent Ising model decomposition [50, 209].
@@
QNN is another representative application without a target unitary matrix. Unlike VQE, which evolves a quantum state according to a given Hamiltonian to obtain the eigenvector corresponding to the minimum eigenvalue, QNN algorithms employ supervised learning. The inspiration comes from classical neural network design methods [105, 65], utilizing layered quantum gate structures to emulate classical neural networks. QNNs are trained to fit training set data, with the loss function typically defined as the prediction accuracy on the test set. For instance, quantum convolutional neural networks [37] and quantum recurrent neural networks [18] incorporate artificially designed quantum layers, which are implemented through parameterized gates and entanglement gates. Similar to VQE, QNNs employ parameterized quantum circuits, calculating the loss function and updating parameters via classical computers. Consequently, these parameterized quantum circuits can be directly obtained by QAS algorithms through the original loss function. QNNs are hypothesized to potentially offer faster training speeds and higher prediction accuracy compared to classical neural networks [83]. Recent studies [208, 187, 50] have utilized QAS algorithms to design complete circuits for QNNs or components of their layers, aiming to achieve improved training effects and modify certain fixed ansatz templates such as the Hardware Efficient Ansatz (HEA).
@@
Beyond these applications, there are other potential applications without target unitary matrices, such as the design of quantum error correction codes and quantum nonlinear units. Quantum error correction codes are crucial in realizing fault-tolerant quantum computing by encoding quantum information to resist environmental noise [29]. The essence of quantum error correction codes involves encoding the original quantum state into a larger state through specific encoding methods to enhance information protection. Consequently, automated circuit synthesis algorithms can be introduced in the encoding and decoding processes of quantum error correction codes. Current research in this domain encompasses heuristic methods [160], machine learning approaches [33, 37], and reinforcement learning techniques [140, 205]. Similarly, the design of quantum nonlinear units is also a potential application area. Since quantum circuits contain only linear unitary transformations, it is difficult for quantum machine learning to introduce nonlinearity as conveniently as in classical machine learning methods [167]. Some recent studies have attempted to use linear oracles to approximate the results of nonlinear operations [159], thus achieving the effect of nonlinear layers to some extent. This makes the design of nonlinear units another potential application area for QAS algorithms.
3.3 Quantum Architecture Search Methods

### v2

@@ II-B Directed Acyclic Graph
To better extract information from quantum circuits, we can use Directed Acyclic Graphs (DAGs) to represent quantum gate circuits and apply them to tasks, e.g., quantum architecture search [57, 58] and quantum compiling [59, 60]. By adding a source node and a sink node at the beginning and end of the quantum circuit, respectively, and treating each quantum gate as a node, we create a directed graph. Adjacent gates on the same qubit are connected by directed edges, and multi-qubit gates introduce additional edges. These directed edges represent the sequence in which the gates are applied. The advantages of using DAGs to represent quantum circuits are: (1) the structure of quantum circuits naturally transforms into DAGs; (2) DAGs can represent the application order of quantum gates and the control qubits of multi-qubit gates, maximizing the advantages of DAGs ; and (3) graph neural networks can effectively extract and compute features of nodes and edges, handling large-scale quantum circuits.
@@
There is a special circuit representation in the form of DAGs called Quantum Multiple-Valued Decision Diagram (QMDD) [61]. As shown in Fig. 2, starting from the $q_{2}$ level, i.e. a $2\times 2$ matrix, we build the diagram layer by layer according to the transformation matrix. Each $2\times 2$ matrix consists of four elements. There are four types of $q_{2}$ level matrices in the circuit, so this level has four nodes. The $q_{1}$ level above comprises a $4\times 4$ matrix. Each $q_{1}$ level matrix is composed of four $q_{2}$ level matrices, connecting each $q_{1}$ level node to four $q_{2}$ level nodes. The final $q_{0}$ level $8\times 8$ unitary matrix is the target unitary matrix we aim to build. The completed QMDD diagram is shown at the bottom of the QMDD module.
II-C Circuit Polynomial
@@
Based on the representation of the quantum gate model, we can further characterize quantum circuits using algebraic methods, e.g. linear Boolean functions or phase polynomials [62, 63, 64, 65, 66]. These polynomial representation algorithms focus more on the quantum state rather than the specific quantum gates acting on the state. An arbitrary single-qubit quantum gate can be seen as inducing the following transformations on the quantum state:
[…]
$p(x_{1},x_{2},x_{3})$ $=\theta_{1}(x_{1})+\theta_{2}(x_{1}\oplus x_{2})$ (2)
$+\theta_{3}(x_{1}\oplus x_{2}\oplus x_{3})+\theta_{4}(x_{2}\oplus x_{3}).$
@@
Since the coefficients in the phase function are all rotation angles limited to a range of $2\pi$, we can take the modulo of the superimposed angle. This allows the phase function to serve as a criterion for determining whether two circuits are equivalent. Compared to directly comparing two unitary matrices, using phase polynomials enables the equivalence assessment of larger-scale circuits. For instance, [67] optimized a 72-qubit quantum circuit based on phase polynomial, a task that cannot be accomplished by merely using unitary matrices.
II-D Tensor Networks
@@
A tensor [68] can be regarded as a multidimensional vector, where scalars, vectors, and matrices can be considered as 0-, 1-, and 2-dimensional tensors, respectively [69]. [11] introduced the graphical representation of tensors into quantum computing, leading to the quantum gate circuits we discussed earlier. Therefore, the quantum gate model is indeed a special form of tensor networks. We will briefly introduce this more fundamental representation, with more detailed content about the tensor network available in [15]. Tensor networks can connect different tensors through indices, an operation called tensor network contraction. Some classic tensor network structures include Matrix Product State (MPS), Projected Entangled Pair States (PEPS), Tree Tensor Network (TTN), and Multi-Scale Entanglement Renormalization Ansatz (MERA) – see Fig. 2.
@@
Since matrix multiplication can be viewed as the contraction of two second-order tensors based on the same index, quantum circuits can naturally be simulated and optimized using tensor networks [70]. The main optimization method is to reduce the cost of contractions, but finding an optimal tensor network contraction sequence is NP-hard [71]. This problem can be further modeled as finding an optimal tree decomposition of the tensor network graph structure [72]. The tree pruning method proposed in the paper can decompose the original graph into multiple parallel subgraphs with the minimum tree depth. [73] proposed index slicing, which divides the entire contraction into sub-tasks that can be executed in parallel. Specifically, a greedy algorithm selects indices that lead to the greatest reduction in complexity, accompanied by local reordering of the contraction tree. Additionally, pre-computing small tensors on the CPU and reordering matrix multiplication on the GPU can reduce the actual running cost. Contraction tasks can be divided into several tensor transpositions and one matrix multiplication [74]. [75] constructed a task graph based on the dependencies of the aforementioned transpositions and multiplications, allowing different transpositions to be parallelized. Furthermore, by naming to ensure that shared operations between sliced contractions are only calculated once, costs are significantly reduced.
@@
Another research direction is to optimize performance by simulating various tensor network structures when designing quantum circuits. These circuits are often referred to as Tensor Network Quantum Circuits (TNQC). Generative and adversarial tasks are completed using quantum circuits based on MPS and TTN [76]. Since TNQC allows us to simulate certain quantum circuits efficiently, we can use this method to save physical qubits and reduce the impact of real quantum computing noise. Quantum circuits based on MPS and MERA are also analyzed in [77], finding that TNQC can use fewer quantum gate parameters to achieve the same accuracy, further validating the practicality of TNQC.
II-E ZX Diagrams
@@
ZX diagrams are a method for representing quantum circuits based on ZX calculus, proposed in [14]. ZX diagrams are also a special form of tensor networks that can represent any linear mapping between qubits [78]. They use spiders and the lines connecting spiders to represent a quantum circuit. Typically, ZX diagrams include two types of spiders: green Z spiders and red X spiders, as shown in Fig. 2 [79]. When the angles are $\pi$, these two types of spiders correspond to the Pauli-Z and Pauli-X gates, respectively.
@@
Spiders define the basic linear operations in any dimension, where the lines on the left side of a spider represent the inputs of the current linear operation, and the lines on the right side represent the outputs. Each spider’s output can further connect to the inputs of other spiders, constructing a large network. ZX diagrams follow a set of fusion rules, which allow for obtaining many equivalent representations, simplifying the connections of complex spiders. Since 2019, much work has been done to optimize quantum circuits based on ZX diagrams. For instance, [80, 81, 82, 83] studied the optimization of the number of spiders in ZX diagrams; [84, 26, 85, 86] researched the optimization of T gate counts; and papers like [87, 88] introduced heuristic algorithms for the contraction and optimization of ZX diagrams. By optimizing the ZX diagram representation of quantum circuits, we can obtain more streamlined quantum circuits. When these ZX diagrams are mapped back to gate circuits, we can get circuits that meet the requirements. Compared to directly optimizing quantum gate circuits, the optimization methods based on ZX diagrams can introduce different characteristics, making them a relatively common method for representing quantum circuits.
III Quantum Logic Circuit Synthesis
@@
In this section, we address the challenge of quantum logic circuit synthesis. Running a quantum algorithm on a quantum processor necessitates translating the algorithm’s unitary transformations into quantum circuits. Typically, these circuits are designed manually [89] or through decomposition algorithms [90, 91, 92]. This process demands substantial human labor and expertise, often resulting in circuits that are particularly deep. In the current NISQ era, various constraints must be considered, including the accuracy of single and double qubit gates, qubit decoherence time, and the topological structure of physical qubits. Constructing a quantum circuit that achieves the highest possible accuracy while adhering to these constraints is a pressing challenge. Therefore, this section delves into methods for automatically generating quantum circuits that meet these requirements. This automated synthesis of quantum logic circuits is also referred to as the Quantum Architecture Search (QAS) problem in literature [93, 19, 94].
III-A Problem Definition
@@
First, we provide the mathematical definition of the QAS problem. The QAS problem originates from the Neural Architecture Search (NAS) problem, which automates the synthesis and optimization of neural network structures and parameters to replace human expert design. NAS has been widely applied in the field of computer vision, e.g. image classification [95] and object detection [96], and has achieved success on multiple datasets. Similar to the NAS setting, when defining the QAS problem, we first determine the set of candidate gates and the search space. Assuming we need to search for a circuit with $n$ qubits and $m$ layers, define the set of candidate gates as $\mathcal{G}$, and the size of the candidate set as $k=|\mathcal{G}|$, then the search space of the QAS problem is of the size $k^{n\times m}$. The candidate gate set $\mathcal{G}$ can include single-qubit, two-qubit, and parameterized quantum gates. We denote the selected candidate gates as matrix $\mathbf{M}\in\mathcal{G}^{n\times m}$, define $\mathbf{U}_{ij}=\sigma(\mathbf{M}_{ij})$, where $\sigma$ maps a quantum gate on the $i$-th qubit and $j$-th layer to a $2^{n}\times 2^{n}$ unitary matrix with all the irrelevant qubits as identity. The unitary matrix corresponding to the searched circuit is:
@@
For different applications, the QAS problem can be further divided into two versions. The first searches the circuit with a target unitary matrix $\mathbf{U}$ and the corresponding objective is:
[…]
@@
III-B Applications of Quantum Architecture Search
III-B1 Quantum Architecture Search with Unitary Matrix
The most important application for QAS with unitary matrix is the circuit implementation of oracles. Here, an oracle refers to a black-box unitary matrix often seen in quantum algorithms. We may not know its specific implementation yet, but we understand the unitary operation corresponding to this oracle. Quantum oracle is of fundamental significance in quantum computing and quantum information, but implementing an arbitrary quantum oracle requires exponentially many gates [48]. Taking the oracle $\mathcal{O}$ in Grover’s algorithm [2] as an example, which can be represented as a unitary operation:
@@ III-B1 Quantum Architecture Search with Unitary Matrix
[…]
@@
Researchers typically manually design these oracles using discrete quantum gates (i.e., quantum gates without parameters) [16, 17]. However, with the proposal of AI-empowered search methods, parameterized quantum circuits have also been used to search for implementations of these oracles more efficiently. Search algorithms are also used for state preparation and operator synthesis [97, 98], evolve quantum oracles [99], as well as design quantum adders [100, 101], which is further applied to quantum autoencoders [102].
III-B2 Quantum Architecture Search without Unitary Matrix
@@
In practical applications, situations without a target unitary matrix are more common, especially in quantum machine learning problems e.g. Quantum Neural Networks (QNN) or Variational Quantum Eigensolvers (VQE). Theoretically, if there are enough linearly independent input-output pairs (quantum initial states and corresponding final states), the unitary matrix corresponding to the transformation circuit is mathematically solvable, making this situation equivalent to the case with a target unitary matrix. However, the search process becomes more challenging when there is a lack of linearly independent input-output pairs.
@@
The VQE algorithm [103] is a notable application without a target unitary matrix, with extensive utility in quantum chemistry, combinatorial optimization, and quantum simulation. Its core functionality involves optimizing parameterized quantum circuits through classical optimizers to determine the minimum eigenvalue and corresponding eigenvector of a given Hamiltonian. VQE has garnered significant attention in the NISQ era due to its relatively shallow parameterized quantum circuits and adaptability to noisy environments on quantum processors. When utilizing VQE to solve quantum chemistry problems as the ground state energy estimation problem, the process typically involves discretizing electron distributions into orbitals through second quantization [104], mapping these orbitals to qubits using algorithms e.g. the Jordan-Wigner transformation [105], and employing unitary coupled cluster theory [106] to generate single and double excitation ansatze for quantum state evolution. This procedure necessitates the derivation of distinct quantum circuits for each molecule or Hamiltonian, requiring both quantum chemistry expertise and specialized algorithms. Moreover, the resulting ansatze are often deeper [107] due to the preservation of specific quantum chemical symmetries, rendering them challenging to implement on current NISQ hardware.
Consequently, researchers have explored QAS methods to automatically design VQE ansatz, aiming to reduce design complexity and generate circuits that satisfy given constraints (e.g., circuit depth, two-qubit gate count, or physical qubit connectivity). The search process also utilizes Hamiltonian expectation as the loss function. Existing research has applied QAS methods to enhance unitary coupled cluster ansatze [18, 108] or directly search for VQE circuits without prior ansatz knowledge [109, 110, 20]. Similarly, in combinatorial optimization problems, e.g. the maximum cut problem solved using the Quantum Approximate Optimization Algorithm (QAOA)[111], studies have proposed methods to optimize the original QAOA circuit[112, 113] or design related circuits that circumvent Ising model decomposition [114, 93].
@@
QNNs represent another application domain that does not require a specific target unitary matrix. Unlike VQE, which evolves a quantum state according to a given Hamiltonian, QNNs are a type of supervised learning algorithm. Inspired by classical neural network architectures [115, 116], they employ layered quantum gate structures to emulate their classical counterparts. QNNs are trained to fit a dataset by minimizing a loss function, which is typically defined based on prediction accuracy during training. For instance, quantum convolutional neural networks [22] and quantum recurrent neural networks [117] incorporate manually designed quantum layers, implemented using parameterized and entanglement gates. Similar to VQE, QNNs use parameterized quantum circuits, and their parameters are updated by a classical optimizer based on the calculated loss function. Consequently, these parameterized quantum circuits can be directly designed using QAS algorithms with the same original loss function. It is hypothesized that QNNs may offer advantages e.g. faster training and higher prediction accuracy compared to their classical counterparts [118]. Recent studies [119, 110, 114] have utilized QAS algorithms to either design complete QNN circuits or optimize specific layers, aiming to achieve improved training outcomes and modify predefined fixed ansatz templates like the Hardware Efficient Ansatz (HEA).
@@
Beyond these applications, there are other potential applications without target unitary matrices, e.g. the design of quantum error correction codes and quantum nonlinear units. Quantum error correction codes are crucial in realizing fault-tolerant quantum computing by encoding quantum information to resist environmental noise [120]. The essence of quantum error correction codes involves encoding the original quantum state into a larger state through specific encoding methods to enhance information protection. Consequently, automated circuit synthesis algorithms can be introduced in the encoding and decoding processes of quantum error correction codes. Current research in this domain encompasses heuristic methods [121], machine learning approaches [21, 22], and reinforcement learning techniques [122, 123]. Similarly, the design of quantum nonlinear units is also a potential application area. Since quantum circuits contain only linear unitary transformations, it is difficult for quantum machine learning to introduce nonlinearity as conveniently as in classical machine learning methods [124]. Some recent studies have attempted to use linear oracles to approximate the results of nonlinear operations [23, 24], thus achieving the effect of nonlinear layers to some extent. This makes the design of nonlinear units another potential application area for QAS algorithms.
III-C Quantum Architecture Search Methods